OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sin(7pi/8)

OpenStudy (solomonzelman):

http://www.purplemath.com/modules/idents.htm

OpenStudy (solomonzelman):

` π=180 `

OpenStudy (solomonzelman):

7×22.5 Or (1/2)×7×45

OpenStudy (solomonzelman):

Go for it. Use your ½∠ sine formula

OpenStudy (anonymous):

what's x?

OpenStudy (solomonzelman):

In your case it would be 315 (and that x is over 2 )

OpenStudy (solomonzelman):

I personally prefer ` sin(a+b) ` though -:( ... anyway ,

OpenStudy (anonymous):

how did you get 315?

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

@ByteMe yes, more help would be GREATLY appreciated!

OpenStudy (anonymous):

@amistre64 @ParthKohli @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Use the identity \[\Large \sin\left(\frac{\theta}{2}\right) = \pm\sqrt{\frac{1-\cos(\theta)}{2}}\]

jimthompson5910 (jim_thompson5910):

In your case, \[\Large \theta = \frac{7\pi}{4}\] because half of this angle is \[\Large \frac{\theta}{2} = \frac{1}{2}*\frac{7\pi}{4}\] \[\Large \frac{\theta}{2} = \frac{7\pi}{8}\]

OpenStudy (anonymous):

so theta is 7pi/4?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so I have sin((7pi/4)/2)=+-sqrt((1-cos(7pi/4)/2) left?

jimthompson5910 (jim_thompson5910):

keep going use the unit circle to evaluate cos(7pi/4)

jimthompson5910 (jim_thompson5910):

and then simplify

OpenStudy (anonymous):

sqrt(2)/2?

OpenStudy (anonymous):

sin((7pi/4)/2)=+-sqrt((1-(sqrt(2)/2)/2)?

jimthompson5910 (jim_thompson5910):

keep going and simplify

OpenStudy (anonymous):

can I simplify sin(7pi/4)?

jimthompson5910 (jim_thompson5910):

no don't worry about the left side right now

OpenStudy (anonymous):

would I distribute the - to the sqrt(2)/2?

jimthompson5910 (jim_thompson5910):

inside the square root, you can multiply every term by 2 to get \[\Large \sin\left(\frac{\frac{7\pi}{4}}{2}\right) = \pm\sqrt{\frac{1-\frac{\sqrt{2}}{2}}{2}}\] \[\Large \sin\left(\frac{\frac{7\pi}{4}}{2}\right) = \pm\sqrt{\frac{2-\sqrt{2}}{4}}\] What's next?

OpenStudy (anonymous):

getting rid of the denominator (4)?

jimthompson5910 (jim_thompson5910):

You can break up the square root

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{\frac{A}{B}} = \frac{\sqrt{A}}{\sqrt{B}}\]

OpenStudy (anonymous):

so it'd be +-(2-sqrt(2)) / sqrt4

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

+-sqrt4?

jimthompson5910 (jim_thompson5910):

\[\Large \sin\left(\frac{\frac{7\pi}{4}}{2}\right) = \pm\sqrt{\frac{2-\sqrt{2}}{4}}\] \[\Large \sin\left(\frac{\frac{7\pi}{4}}{2}\right) = \pm\frac{\sqrt{2-\sqrt{2}}}{\sqrt{4}}\]

OpenStudy (anonymous):

\[+-\frac{ \sqrt{2-\sqrt{2}} }{ 2 }\]

jimthompson5910 (jim_thompson5910):

So that means \[\Large \sin\left(\frac{7\pi}{8}\right) = \pm\frac{\sqrt{2-\sqrt{2}}}{2}\] -------------------------------------------------------------------- Since sin(7pi/8) = 0.38268343236 we know that the final answer is positive, so, \[\Large \sin\left(\frac{7\pi}{8}\right) = \frac{\sqrt{2-\sqrt{2}}}{2}\]